The sum of the ages of Musa and his friend is 50 years.If ten years ago,the age of Musa was twice that of his friend.Find their present Ages.
Answer by jorel555(1290)

Let m be Musa,n be his friend. Then
m+n=50
n=50-m;
n-10=1/2(m-10)
80-2m-=m-10
3m=90
m=30
n=20!!!!!!!!!!!!!